OpenStudy (anonymous):

what is the inverse of y = -2x - 4 ? HELP PLEASE

OpenStudy (anonymous):

you just need to replace x with y and y with x.

OpenStudy (anonymous):

Y=-2x-4 2x+4=-y 2y+4=-x x+4=-2y -(x+4)/2=y -x-4/2=y is your answer. y=(-x-4)/2